the perimeter of a certain soccer field is 310m. the lenght is 65 more than the width. what are the dimension?
Answer by benazir.sj@gmail.com(109)

let width =x
length =x+65
given perimeter =310
perimeter =2(l+w)
310=2(x+65+x)
310=2(2x+65)
310=4x+130
310-130=4x
180=4x
x=180/4
x=45
therefore;
width =45
and length =x+65
=45+65
=110